OpenStudy (anonymous):

Consider a square with side 54.1 cm. Calculate the electric field at the center if one corner is occupied by a +41.9 µC charge and the other three are occupied by -27.3 µC charges.

OpenStudy (theeric):

Hi! Do you have any idea where to start? Have you attempted a solution?

OpenStudy (anonymous):

yea i did but i cant seem to get the right answer

OpenStudy (theeric):

Let me see what I get, then...

OpenStudy (theeric):

Ar you using Coulomb's constant as \(8.987\times 10^9\left[\dfrac{\rm{N\bullet m^2}}{\rm{C^2}}\right]\)?

OpenStudy (anonymous):

yea E=k*q/r^2

OpenStudy (anonymous):

where r = sqrt2*54.1

OpenStudy (theeric):

And you used \(r=\sqrt{54.1^2+54.1^2\ }\)?

OpenStudy (theeric):

\(\implies r^2=2(54.1)^2=5854\ \rm m^2\)

OpenStudy (theeric):